OpenStudy (anonymous):

Marquis has a gross pay of $816. By how much will his gross pay be reduced if Marquis has the following items withheld? federal tax of $92 Social Security tax that is 6.2% of his gross pay Medicare tax that is 1.45% of his gross pay state tax that is 22% of his federal tax a. $82.66 b. $129.13 c. $164.01 d. $174.66

OpenStudy (anonymous):

please help! will give medal!

OpenStudy (anonymous):

816*0.062 = 50.592 816*0.0145 = 11.832 92*0.22 = 20.24 ------------------------ 82.664 (other taxes) + the 92$ tax = 174.66